Previous | Calendar |
July |
2007 |
Next | |||||||||
---|---|---|---|---|---|---|---|---|---|---|---|---|---|
Sunday | Monday | Tuesday | Wednesday | Thursday | Friday | Saturday | |||||||
1JD182 (clear/hot)Site Visit (SpB): Details , Photos (7) more work on HYDRA (using thinkpad) |
2JD183 |
3JD184Site visit (JB,SpB,LL,RM,Aaron): Details , Photos (3) Moved hydra to creek location Downloaded ALL logger data |
4JD185Independence Day |
5JD186Site visit (JB,SpB,NT,SL,Ramesh): Details , Photos (15) JB swapped 15 flasks Put "forestcam" on cu tower |
6JD187from 11:36-16:55 MDT, archive process on russter2 died. |
7JD188 |
|||||||
8JD189 |
9JD190Site visit (JB,SpB): Details , Photos (8) from 9:50-10:08 MST, Changed profiler (li-6251) dessicant; re-zeroed at 10:03 MST Got webcam back on tower and working deployed 2nd li-cor PAR sensor (s/n Q22845?) w/ shading band at top of tower (PAR2.up) Removed old Fortress UPS (+ disconnected outlets in breaker box) |
10JD191Site visit (RM, SD, SO, + others): Details Site visit by Scott Denning group (CSU) Tram run (with video) |
11JD192Site visit (JB, SpB, RKM, EB, SL): Details , Photos (12) Dismantled former east-robot tower, added 2 sections to the NCAR "central" tower (from NIWOT02) |
12JD193Site Visit (JB, MP, JH, DK, SLS, Ramesh, LL, SL, NT): Details , Photos (6) JB swapped 15 flasks JB took forest phenol photos |
13JD194Site visit (JB, NT, AT): Details JB guyed 3 corners of Maggie's tower JB swapped N2 cylinders at the base of the CU tower new fast ozone target (AT) |
14JD195Site Visit (SpB): Details , Photos (9) from 12:00-13:00 MST, periods of hail/rain. |
|||||||
15JD196 |
16JD197 |
17JD198Site Visit (JB,SpB,LL,RM,JH,MP,NT,SLS,DK): Details , Photos (14) replaced the 16m propvane's cable but this did not prevent the signal from going negative Moved webcam to ~17m at 9:23 MST, disconnected 2nd REBS Rnet sensor from CNR logger and moved to mini-tower/UCB logger. (note, the fan is not connected). at UCB logger started running ucb_08.csi |
18JD199 |
19JD200JB has dentist appointment at 9:45 a.m. h2o.ec clipped during day (at ~1050 mV). |
20JD201Site Visit (JB,LL,Ramesh,SpB): Details , Photos (14) JB and SpB moved N2 and ~1m "test" inlets over to hydra JB swapped 15 flasks from 10:14-10:16 MST re-zeroed LI-6262, h2o.ec set to -30mV to try and keep in A-to-D range. downloaded data from RAD, PROF, soil, and Tcs removed 16m propvane for further trouble shooting (with Kurt at the Marr Lab) |
21JD202 |
|||||||
22JD203 |
23JD204 |
24JD205 |
25JD206 (cool/wet)Site Visit (JB,LL,RKM,SpB,RM,JS,AT): Details , Photos (5) at 9:40 MST swapped mag. perchlorate trap in Li-6251, during blow test found a BIG LEAK!! (returned at 14:00 MST and found a loose ferrule on left side of nupro filter, replaced ferrule and then it seemed ok) added shadow band back to "PAR2.up" (still needs adjustment) from ~11:00-12:00 MST, heavy hail/rain. |
26JD207Site Visit (JB,SL,JH,SLS): Details JB swapped 15 flasks WSU starts SF6 measurements (or maybe previous night?) |
27JD208Site Visit (JB,SpB,RKM,Ramesh,RM): Details RKM placed the "CU tower" hydra line at the lowest inlet height on the mini tower (blow test started at 12:20 MST, took about 5 minutes to reach LI-7000) JB and SpB calibrated the LI-COR 7000 at hydra at HYDRA: 16 inlets connected. at 11:50 MST, start collecting co2 data. lineflow set to "40" from ~14:00-15:00 MST, intense rain shower. |
28JD209
|
29JD210
|
30JD211Site Visit (JB,DK,SL,NT): Details SL and NT started diurnal soil respiration measurements at dawn; JB and DK helped at HYDRA: at 9:45 MST, dropped dry ice into HYDRA box (JB) Trade in jeep and pick up suburban
|
31JD212
|
1JD213Site Visit (JB, RKM, SpB, RaM, RM): Details , Photos (9) JB replaced Jia's electrical board at west 90 at ~12:00 MST, cleaned krypton hygrometer optics downloaded from soil, Tcs, RAD, PROF, MET, NCAR Tcs from 12:15-12:30 MST, adjusted shading band so PAR2.up sensor was shaded WSU leaves (to return next week)
|
2JD214Site Visit (JB, SL): Details Brought 3 N2 cylinders to the trailer (CC125575, CC193045, CC198701) Added salt to both grounding pipes closed main valves on both hydra gas cylinders, and powered down hydra swapped 15 flasks
|
3JD215
|
4JD216 |